A pair of boots, originally priced at £60, were reduced to £36 in a sale. What was the percentage reduction in the price of the boots? Thank you so much :)
Answer by waynest(281)

60 - 36 = 24
24/60 = 0.4
100 * 0.4 = 40%
check:
60 - (60 * 0.4) = 36
60 - (24) = 36
60 - 24 = 36
36 = 36